OpenStudy (anonymous):

Find the gradient of the normal to the curve y=x^2(x^2-9/x) at the point (2,-2) Just to double check my answer @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the equation is this right? \[\Large y = x^2\left(\frac{x^2-9}{x}\right)\]

OpenStudy (anonymous):

no it should be\[y=x^2(x^2-\frac{ 9 }{ x })\]

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

yw

jimthompson5910 (jim_thompson5910):

what answer did you get?

OpenStudy (anonymous):

\[-\frac{ 1 }{ 23 }\]

jimthompson5910 (jim_thompson5910):

I'm getting 23 as the slope of the tangent line at (2,-2) so yes, the slope of the normal line is -1/23

OpenStudy (anonymous):

Thank you @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

you're welcome